# Arnaud Pannatier 1122    Evann Courdier 1122    François Fleuret 33

### v1

@@ Abstract
Autoregressive models, such as the GPT family, use a fixed order, usually left-to-right, to generate sequences. However, this is not a necessity. In this paper, we challenge this assumption and show that by simply adding a positional encoding for the output, this order can be modulated on-the-fly per-sample which offers key advantageous properties. It allows for the sampling of and conditioning on arbitrary subsets of tokens, and it also allows sampling in one shot multiple tokens dynamically according to a rejection strategy, leading to a sub-linear number of model evaluations. We evaluate our method across various domains, including language modeling, path-solving, and aircraft vertical rate prediction, decreasing the number of steps required for generation by an order of magnitude.
@@ 2.3 Conditional Probabilities and Infilling
Figure 2: Left. We can infill the sequence by conditioning on the known part (black points). Right. We can also have estimates of the density at any point of the sequence.
@@ 2.4 Token-based Rejection Sampling
     In parallel sample $T-t$ $r_{t}\sim U[0,1]$ from a uniform distribution
     In parallel, compute the acceptance decision $a_{i}=r_{i}<\min(1,\frac{p(\tilde{x}_{i}|{\mathbb{X}})}{q(\tilde{x}_{i}|{\mathbb{X}},\tilde{x}_{\sigma_{<i}})})$ for every order
     Select the order that accepts the most tokens before seeing a first rejection
     Keep that order and add the $a$ accepted tokens before the first rejection to the prompt
@@ 2.5 Other orders
Our double positional encoding scheme allows for training and evaluating models in any order. Using a randomized order during training allows conditional density estimation, infilling, and burst sampling at inference time. However the double positional encoding scheme allows any order to be used, and it can be used to train models in a deterministic order that is not left-to-right. As an example, we use a deterministic ‘fractal‘ order to see how it compares to a random or left-to-right order. This order starts in the middle of the sequence then recursively goes to the first quarter and three-quarters of the sequence, and goes on recursively until all the positions have been visited. Such an order is fully deterministic, yet we make the hypothesis that this order leads to more difficult training for the model as it cannot rely on the locality of the information. We present the results in Section 3.5. Note that under perfect models, the order of modeling and decoding should not matter because of the chain rules of probability. We give more details about it in Section 0.A.1 of the supplementary material.
@@ 3.1 General performance
Additionally to these tasks, we created a synthetic benchmark for evaluating our burst sampling algorithm.
@@ 3.3 Curriculum Learning
3.4 Open Text Generation: t-SNE of generated sequence
To get a qualitative sense of the generated text by the different methods, We generate 3000 sequences of 1024 tokens with each method, embed each sequence using an embedding model, and then project the embeddings to 2D using t-SNE. We present the results in Figure 4. We used Open-AI text-embedding-3-small [12] to embed the generated sequences into a single 1536 vector embedding. We represent as green embeddings of sequences of the validation set, used as reference. We compute the t-SNE using the whole 15’000 embeddings and then plot each method (blue) and the other considered method (small gray dots). We first see that embeddings of GPT, $\sigma$-GPT, $\sigma$-GPT with burst sampling, and diffusion are spread over the whole space, showing that the model can generate sequences that are coherent with the validation set.
@@ 3.4 Open Text Generation: t-SNE of generated sequence
Figure 4: 2D t-SNE of text-small-3-embeddings of 3000 sequences generated by each method. We compute the t-SNE of all the embeddings together, and then we display in each graph the embeddings of the validation set (green), the embeddings of the corresponding method (blue), and the embeddings of the other methods (gray). We see that the embeddings of the generated sequences have the same overall distribution compared to validation sets, which seems to indicate that $GPT$, $\sigma$-GPT, $\sigma$-GPT with burst sampling, and diffusion models can generate sequences of similar quality.
3.5 Fractal order
@@ 4 Related works
4.1 Shuffling in language models
In exploring the landscape of shuffling used by autoregressive models, one finds noteworthy precedence in the work of [18]. XLNet, which uses randomization of the sequence order as a pretraining task for sentence encoding in the context of natural language understanding. Our approach is similar with regards to the shuffling of the sequence, and the shuffled language modeling but differs in its implementation: we use double positional encoding and a regular causal mask instead of relying on two streams and the modification of the attention matrix. The objective of the two models differs as well, XLNET is used to encode information similarly to BERT while our approach is generative. A very recent approach by [6], trains a transformer both on a left-to-right and right-to-left order which solves a classic problem of the transformer model to understand the context of the token in both directions. Our approach by default trains the model in any order, and we can consider that it sees an exponential number of different sequences during training. However, our section on memoization shows that this does not serve as a good data augmentation technique and that the model seems to memorize the data more in that case.
4.2 Burst sampling scheme
Other works are trying to solve the problem of the linear time required by autoregression using burst sampling. Maskgit [3], for example, uses a BERT-like masked language model (MLM) and uses a custom decoding scheme, which samples multiple tokens at the same time to generate an output. The number of tokens generated at each pass is fixed by the masking schedule and the model utilizes a confidence-driven decoding scheme to choose which tokens to predict. In a distinct approach, [10] introduced a method that leverages an auxiliary model to guide the generation process. Alternatively, the approach outlined by [9] primarily focuses on generating preliminary drafts of an image, which are then progressively enhanced in subsequent iterations. Following work in video generation [2, 16], are leveraging a MaskGit [3] like approach for generation as autoregressive generation of video would be too costly. Our rejection sampling scheme allows as well to generate the sequence by burst but in contrast to another scheme, the number of tokens accepted is dynamic and depends on the data being modelled, which allows faster generation when the underlying data distribution is simple.

### v2

@@ Abstract
Autoregressive models, such as the GPT family, use a fixed order, usually left-to-right, to generate sequences. However, this is not a necessity. In this paper, we challenge this assumption and show that by simply adding a positional encoding for the output, this order can be modulated on-the-fly per-sample which offers key advantageous properties. It allows for the sampling of and conditioning on arbitrary subsets of tokens, and it also allows sampling in one shot multiple tokens dynamically according to a rejection strategy, leading to a sub-linear number of model evaluations. We evaluate our method across various domains, including language modeling, path-solving, and aircraft vertical rate prediction, decreasing the number of steps required for generation by an order of magnitude  [^1] The code of this work is available at https://github.com/idiap/sigma-gpt.
@@ 2.3 Conditional Probabilities and Infilling
Figure 2: (Left.) We can infill the sequence by conditioning on the known part (black points). (Right.) We can also have estimates of the density at any point of the sequence.
@@ 2.4 Token-based Rejection Sampling
     In parallel sample $T-t$ variables $u_{i}\sim U[0,1],\forall i\in t,\dots,T$ from a uniform distribution.
     In parallel, compute the acceptance decision $a_{i}=u_{i}<\min\left(1,\frac{q\left(\tilde{x}_{i}|{\mathbb{X}},\tilde{x}_{\sigma_{<i}}\right)}{p(\tilde{x}_{i}|{\mathbb{X}})}\right)$ for every order.
     Select the order that accepts the most tokens before seeing a first rejection.
     Keep that order and add the $a$ accepted tokens before the first rejection to the prompt.
@@ 2.5 Other Orders
Our double positional encoding scheme allows for training and evaluating models in any order. Using a randomized order during training allows conditional density estimation, infilling, and burst-sampling at inference time. However the double positional encoding scheme allows any order to be used, and it can be used to train models in a deterministic order that is not left-to-right. As an example, we use a deterministic ‘fractal‘ order to see how it compares to a random or left-to-right order. This order starts in the middle of the sequence then recursively goes to the first quarter and three-quarters of the sequence, and goes on recursively until all the positions have been visited. Such an order is fully deterministic, yet we make the hypothesis that this order leads to more difficult training for the model as it cannot rely on the locality of the information. We present the results in Section 3.5. Note that under perfect models, the order of modeling and decoding should not matter because of the chain rules of probability. We give more details about it in Section 0.A.1 of the supplementary material.
@@ 3.1 General performance
Additionally to these tasks, we created a synthetic benchmark for evaluating our burst-sampling algorithm.
@@ 3.3 Curriculum Learning
3.4 Open Text Generation: t-SNE of Generated Sequences
To get a qualitative sense of the generated text by the different methods, We generate 3000 sequences of 1024 tokens with each method, embed each sequence using an embedding model, and then project the embeddings to 2D using t-SNE. We present the results in Figure 4. We used Open-AI text-embedding-3-small [15] to embed the generated sequences into a single 1536 vector embedding. We represent as green embeddings of sequences of the validation set, used as reference. We compute the t-SNE using the whole 15’000 embeddings and then plot each method (blue) and the other considered method (small gray dots). We first see that embeddings of GPT, $\sigma$-GPT, $\sigma$-GPT with burst-sampling, and diffusion are spread over the whole space, showing that the model can generate sequences that are coherent with the validation set.
@@ 3.4 Open Text Generation: t-SNE of Generated Sequences
Figure 4: 2D t-SNE of text-small-3-embeddings of 3000 sequences generated by each method. We compute the t-SNE of all the embeddings together, and then we display in each graph the embeddings of the validation set (green), the embeddings of the corresponding method (blue), and the embeddings of the other methods (gray). We see that the embeddings of the generated sequences have the same overall distribution compared to validation sets, which seems to indicate that $GPT$, $\sigma$-GPT, $\sigma$-GPT with burst-sampling, and diffusion models can generate sequences of similar quality.
3.5 Training and Generating in Fractal Order
@@ 4 Related works
Shuffling in Language Models:
The objective of $\sigma$-GPT is to model shuffled sequences. XLNet [23] uses a similar objective and shuffling of the sequence order as a pretraining task for sentence encoding in the context of natural language understanding. While both approaches are modeling shuffled sequences, they still differ in their implementations: we use a double positional encoding and a regular causal mask instead of masking based on two streams and the modification of the attention matrix. The two models differ as well in their applications, XLNet is used to encode information similarly to BERT while our approach is generative. A recent approach by [6], trains a transformer both on a left-to-right and right-to-left order to solve a classic problem of the transformer model to understand tokens relations in both directions. In a similar setting, it has been shown that text corpora have a preferred left-to-right order of generation [16] and that training in reverse order can lead to a decrease in performance. This is a possible explanation of why training $\sigma$-GPT on text generation needed extra care compared to other tasks. In vision, the raster-scan order is mainly used to unfold 2D patches into sequences. However, a recent work by [9] showed that using a set of predefined orders of unfolding the images can improve performance. They pass the order information not by using a double positional encoding, but by a fixed number of beginning-of-sequences tokens.
Burst-Sampling Scheme:
Other works are trying to solve the problem of the linear time required by autoregression using burst-sampling. Maskgit [3], for example, uses a BERT-like Masked Language Model (MLM) and a custom decoding scheme, which samples multiple tokens at the same time to generate an output. The number of tokens generated at each pass is fixed by a masking schedule and a confidence-driven decoding scheme is used to choose which tokens to predict next. Another approach, [12] relies on an auxiliary model to guide the generation process. Alternatively, the approach of [11] focuses on generating preliminary drafts of an image, which are then iteratively improved. Current Video generation methods [2, 21] are leveraging a MaskGit-like approach [3] for generation because autoregressive generation of video frames would be too costly. Our rejection sampling scheme allows to generate the sequence by burst but in contrast to other schemes, the number of tokens accepted is dynamic and depends on the data being modeled. This allows for faster generation when the underlying data distribution is simple.
Discrete Diffusion Models:
Diffusion models are also able to generate sequences in a few steps. We compared our approach with a discrete diffusion baseline [1]. In the original work, discrete diffusion was also used to generate text, however without the possibility of conditional density estimation and infilling. Most of the diffusion approaches do not support conditional density estimation and infilling by default. Consistency Models [18], which be adapted from continuous diffusion models, can be used to infill images but in a continuous case. However, there have been recent discrete text-diffusion models that allow infilling [13, 7].